Use the information about a pair of similar solids to find all possible values of x:
Scale factor is 1:2; Volume ration is x-2:x^2-1

Answer by khwang(438)

For the two similar solids, the scalar factor 1:2 means their ratio in
length. Hence, their ratio in volume should be 1^3:2^3 = 1:8.
Now, we have (x-2):(x^2-1) = 1:8,
So, x^2-1 = 8(x-2), or x^2- 8x + 15 = 0.
Factoring (x-3)(x-5) = 0, we have x=3 or 5.
Kenny